OpenStudy (michaelbp):

anyone know how to do half life? i need help with some problems just need someone to help with all

OpenStudy (3mar):

physics?

OpenStudy (3mar):

How many days?

OpenStudy (michaelbp):

im guessing just one day

OpenStudy (eliesaab):

Half life formula \[ \frac{\log (2)}{k} \]

OpenStudy (michaelbp):

im guessing k is???

OpenStudy (eliesaab):

\[ \frac{\log (2)}{0.1325}=5.2313 \]

OpenStudy (michaelbp):

ok thanks and thats how you do them all?

OpenStudy (eliesaab):

yes

OpenStudy (michaelbp):

ok thanks ;)

OpenStudy (3mar):

Did you get it?

OpenStudy (leenathan):

@michaelbp u need to shave or let ur beard grow out

OpenStudy (leenathan):

and also put your hair up and get muscle like me

OpenStudy (leenathan):

just some advice

OpenStudy (michaelbp):

lol ok thanks